If the point (H,K ) lies on the line 2x+3y=6 , the smallest value of H square+ k square is?

Grade:11

1 Answers

Vikas TU
7 years ago
If the point (H,K ) lies on the line 2x+3y=6 , the smallest value of H square+ k square would be as follows:
put (H,K) in the given eq.
2h + 3k = 6
or
k = (6 – 2h)/3
Now,
let
M = h^2 + k^2
put value of k in M and differentiate M w.r.t h and equate it to zero.
then put that h values in M to get the minum value.